media%2F083%2F083b1c0e-df3d-4457-a765-00
Answer #1

given 1 fluid ounce= 29.573 ml

again 1L= 1000ml= 1000cm^3

now density means per 1 cm^3 volume the water mass is 1gm

now if 1000ml =1000 cm^3

then 1ml = 1cm^3

1gm= 10^-3kg

so 10^-3 Kg water is in 1ml volume

or 1kg water is in 1/10^-3 ml volume = 10^3 ml

again 29.573 ml = 1 fluid ounce

or 1ml = 1/29.573 fluid ounce

or 10^3 ml = 10^3 * 1/29.573 fluid ounce = 33.81fluid ounce